# On the Robustness of the Successive Projection Algorithm

### v1

@@ 6.1 Tightness of SPA2
[16, Theorem 4.1] Under the same assumptions and bounds on the noise as Theorem 1, the index set $\mathcal{J}^{\prime}$ extracted by SPA2 satisfies the following error bound:
$\min_{\pi\in S^{r}}\max_{1\leq k\leq r}\|\|W(:,k)-X(:,\mathcal{J}^{\prime}_{\pi(k)})\|_{2}\leq\mathcal{O}\left(\varepsilon\mathcal{K}(W)\right).$
@@ 6.2 Tightness of MVE-SPA
[18, Theorem 2.9] Let $X=WH+N\in\mathbb{R}^{r\times n}$ where $\operatorname{rk}(W)=r$, $H\in\mathbb{R}^{r\times n}_{+}$ is a separable matrix satisfying $H^{\top}e\leq e$, and
@@
$\min_{\pi\in S^{r}}\max_{1\leq k\leq r}\|W(:,k)-X(:,\mathcal{J}_{\pi(k)})\|_{2}\leq\mathcal{O}\left(\varepsilon\mathcal{K}(W)\right).$
@@
As a consequence, the minimum volume ellipsoid containing all column of $X$ must have the third column on its boundary, and MVE-SPA may identify it after the preconditioning, since all the columns of the preconditioned matrix $\widetilde{X}$ have norm $1$ as it can be seen in Figure 6. The distance between the third column and the first is at least
@@ 7 Translation and Lifting
In this section, we propose a new strategy, combining the lifting procedure with a translation, in order to minimize the conditioning of $W$. In fact, under the condition that $H^{\top}e=e$, translating and shifting does not change the properties of the problem since
@@
In the following, we propose a simple heuristic to tackle this problem. In order to decrease the largest singular value of $W-ve^{\top}$, a heuristic is to minimize the sum of its squared singular values, i.e. $\|W-ve^{\top}\|_{F}^{2}$. The optimal solution of this problem is given by the average of the column of $W$, that is, $v=\frac{1}{r}We$. Note that, if it happens that the best rank-one approximation of $W$ has the form $ue^{\top}$ for some vector $u$ (in other terms, $e/\sqrt{r}$ is the first right singular vector of $W$), then the largest singular value of $W-ue^{\top}$ becomes $\sigma_{2}(W)$ which is the best possible scenario. In practice, this average is typically well approximated by the average of the columns of $X$, given that the data points are evenly distributed in $\operatorname{conv}(W)$. Hence, we suggest to use $v=\frac{1}{n}Xe$. It is also possible to use more advanced strategies, e.g., take $v$ as the average of the columns of $X$ extracted by SPA [1]. However, for simplicity, we consider in this paper only $v=\frac{1}{n}Xe$.
Then, assuming $v=\frac{1}{r}We$, we can characterize the conditioning of $\binom{W-ve^{\top}}{ce^{\top}}$ as a function of $c$ as follows, from which we will propose a value for $c$.
@@
According to (6), assuming $v=We/r$, the value of $c$ minimizing $\mathcal{K}(W_{\ell})$ is in the interval
@@
How can approximate this interval since $W$ is unknown? Since the columns of $X_{t}=X-ve^{\top}$ are within the convex hull of the columns of $W_{t}=W-ve^{\top}$, we assume $\|W_{t}\|_{F}\approx\frac{\sqrt{r}}{\sqrt{n}}\|X_{t}\|_{F}$, and hence we approximate the upper bound of the interval using the following approximation: $\left\|W_{t}\right\|_{F}/\sqrt{r}\approx\left\|X_{t}\right\|_{F}/\sqrt{n}$. For the lower bound, we estimate $\sigma_{r-1}(W_{t})$ with $\min_{j}\|X_{t}(:,j)\|_{2}$. Finally, we propose to use $c$ as the average of the estimated lower and upper bounds,
$c=\frac{1}{2}\left(\min_{i}\|X(:,i)-v\|_{2}/\sqrt{r}+\left\|X-ve^{\top}\right\|_{F}/\sqrt{n}\right).$
@@
We will refer to the variant of SPA that applies SPA on $W_{\ell}$, the translated and lifted data, as translated lifted SPA (TL-SPA).
@@ 8 Numerical experiments
We do not include MVE-SPA as it performs similarly as SPA2, as reported, e.g., in [14, Chapter 7]. The code used to run these experiments is available from https://gitlab.com/ngillis/robustSPA. We do not report computational times as all algorithms are extremely fast, although FAW is about $r$ times slower than SPA, T-SPA and TL-SPA, while SPA2 and TL-SPA2 are about two times slower. (It takes about 10 seconds on a standard laptop to run the experiments described below, running the 6 algorithms on 6120 matrices.)
@@
We use the same synthetic data sets as in  [14, Chapter 7.4.6.1]: $X=WH+N$ where
  • The entries of $W\in\mathbb{R}^{m\times r}$ are generated uniformly at random in the interval $[0,1]$, W = rand(m,r). In this case, $W$ is well-conditioned with high probability. To make $W$ ill-conditioned, we compute the compact SVD of $W=U\Sigma V^{\top}$, then update $W\leftarrow U\Sigma^{\prime}V^{\top}$ where $\Sigma^{\prime}$ is a diagonal matrix whose entries are log-spaced between $10^{-6}$ and 1, hence $\kappa(w)=10^{6}$.
@@
For each experimental settings, we test 51 different noise levels $\delta$ and, for each noise level, we report the average accuracy (which is the proportion of the indices of the columns of $W$ correctly recovered by the corresponding algorithm) over 30 randomly generated matrices. Table 2 reports the robustness of each algorithm in each experimental setting. The robustness is defined as the largest noise level for which the corresponding algorithm recovered exactly all indices for all matrices.
@@
  • For the well-conditioned Dirichlet experiment (Exp. 1), all SPA variants perform similarly and have the same robustness. The main reason is that the conditioning of $W$ and its translated variant, $W_{t}$, are close to each other. This is a similar observation as in [14, Fig 7.6].
@@
  • For the well-conditioned middle point experiment (Exp. 3), there is a clear hierarchy: SPA and T-SPA perform similarly (since the matrix is well-conditioned, translation does not bring much improvements), FAW is more robust than TL-SPA but its accuracy decays more rapidly, SPA2 outperforms FAW (as already noted in [14, Fig 7.7]), while TL-SPA2 and SPA2 perform similarly although TL-SPA2 is more robust (see Table 2).
@@ 9 Conclusion and further work
In this paper, we revisited the robustness to noise of SPA and variants. We improved previous error bounds for the first step of SPA and when $r=2$ (Theorems 3 and 4), and for the first two steps and when $r=3$ for the translated variant of SPA, T-SPA (Theorems 5 and 6). We then proved that, for $r\geq 3$, the original robustness result of SPA [17, Theorem 3] is tight (Theorem 7), an important open question in the literature [14, p. 231], using a well-constructed family of matrices that achieve the worst-case error bound. We also proved the tightness of the error bounds for two preconditioned variants of SPA, namely SPA preconditioned with itself (SPA2, Theorem 9) and SPA preconditioned with minimum-volume ellipsoid (MVE-SPA, Theorem 11). Finally, we proposed a new SPA variant that uses a translation and lifting of the data points, namely TL-SPA, and that allows us to improve the robustness of SPA by reducing the condition number of $W$. We illustrated this on numerical experiments.
@@
Exp. 1: Well-condition Dirichlet experiment
Exp. 2: Ill-condition Dirichlet experiment
Figure 7: Accuracy vs. noise level for Dirichlet experiments.
Exp. 3: Well-condition middle point experiment
Exp. 4: Rank-deficient middle point experiment
Figure 8: Accuracy vs. noise level for the middle point experiments.

### v2

@@ 6.1 Tightness of SPA2
[18, Theorem 4.1] Under the same assumptions and bounds on the noise as Theorem 1, the index set $\mathcal{J}^{\prime}$ extracted by SPA2 satisfies the following error bound:
$\min_{\pi\in S^{r}}\max_{1\leq k\leq r}\|W(:,k)-X(:,\mathcal{J}^{\prime}_{\pi(k)})\|\leq\mathcal{O}\left(\varepsilon\mathcal{K}(W)\right).$
@@ 6.2 Tightness of MVE-SPA
[20, Theorem 2.9] Let $X=WH+N\in\mathbb{R}^{r\times n}$ where $\operatorname{rk}(W)=r$, $H\in\mathbb{R}^{r\times n}_{+}$ is a separable matrix satisfying $H^{\top}e\leq e$, and
@@
$\min_{\pi\in S^{r}}\max_{1\leq k\leq r}\|W(:,k)-X(:,\mathcal{J}_{\pi(k)})\|\leq\mathcal{O}\left(\varepsilon\mathcal{K}(W)\right).$
@@
As a consequence, the minimum volume ellipsoid containing all column s of $X$ must have the third column on its boundary, and MVE-SPA may identify it after the preconditioning, since all the columns of the preconditioned matrix $\widetilde{X}$ have norm $1$ as it can be seen in Figure 6. The distance between the third column and the first is at least
@@ 7 Translation and Lifting
In this section, we propose a new strategy, combining the lifting procedure with a translation, in order to minimize the conditioning of $W$. In fact, under the condition that $H^{\top}e=e$, translating and shifting does not change the solution $H$ and translates and shifts $W$ in the same way since
@@
In the following, we propose a simple heuristic to tackle this problem. In order to decrease the largest singular value of $W-ve^{\top}$, a heuristic is to minimize the sum of its squared singular values, that is, $\|W-ve^{\top}\|_{F}^{2}$. The optimal solution of this problem is given by the average of the column of $W$, that is, $v=\frac{1}{r}We$. Note that, if it happens that the best rank-one approximation of $W$ has the form $ue^{\top}$ for some vector $u$ (in other terms, $e/\sqrt{r}$ is the first right singular vector of $W$), then the largest singular value of $W-ue^{\top}$ becomes $\sigma_{2}(W)$ which coincides with the smallest $\sigma_{1}(W-ze^{\top})$ among all $z$ (by the interlacing property of singular values applied to rank $1$ modifications). In practice, this average is typically well approximated by the average of the columns of $X$, given that the data points are evenly distributed in $\operatorname{conv}(W)$. Hence, we suggest to use $v=\frac{1}{n}Xe$. It is also possible to use more advanced strategies, e.g., take $v$ as the average of the columns of $X$ extracted by SPA [1]. However, for simplicity, we consider in this paper only $v=\frac{1}{n}Xe$. Then, assuming $v=\frac{1}{r}We$, we can characterize the conditioning of $\binom{W-ve^{\top}}{ce^{\top}}$ as a function of $c$ as follows, from which we will propose a value for $c$.
@@
According to (6), assuming $v=We/r$, and denoting by $W_{t}=W-ve^{\top}$ the $t$ranslated version of $W$, and by $W_{\ell}=\begin{pmatrix}W_{t}\\ c\;e^{\top}\end{pmatrix}$ the $\ell$ifted version of $W$, the value of $c$ minimizing $\mathcal{K}(W_{\ell})$ is in the interval
@@
How can approximate this interval since $W$ is unknown? Since the columns of $X_{t}=X-ve^{\top}$ are within the convex hull of the columns of $W_{t}=W-ve^{\top}$, we assume $\|W_{t}\|_{F}\approx\frac{\sqrt{r}}{\sqrt{n}}\|X_{t}\|_{F}$, and hence we approximate the upper bound of the interval using the following approximation: $\left\|W_{t}\right\|_{F}/\sqrt{r}\approx\left\|X_{t}\right\|_{F}/\sqrt{n}$. For the lower bound, we estimate $\sigma_{r-1}(W_{t})$ with $\min_{j}\|X_{t}(:,j)\|$. Finally, we propose to use $c$ as the average of the estimated lower and upper bounds,
$c=\frac{1}{2}\left(\min_{i}\|X(:,i)-v\|/\sqrt{r}+\left\|X-ve^{\top}\right\|_{F}/\sqrt{n}\right).$ (7)
@@
We will refer to the variant of SPA that applies SPA on $W_{\ell}$, the translated and lifted data, as translated lifted SPA (TL-SPA) ; see Algorithm 3.
Algorithm 3 Translated lifted SPA (TL-SPA)
1:  $v=Xe/n$.
2:  $c=\frac{1}{2}\Big{(}\min_{i}\|X(:,i)-v\|/\sqrt{r}+\left\|X-ve^{\top}\right\|_{F}/\sqrt{n}\Big{)}$.
3:  $X_{t\ell}=\begin{pmatrix}X-ve^{\top}\\ c\;e^{\top}\end{pmatrix}$.   % Translate and lift
4:  $\mathcal{J}=$ SPA($X_{t\ell}$,$r$).    % see Algorithm 1
Remark 4 (Implementation of TL-SPA).
As for SPA and T-SPA (see Remarks 1 and 2), TL-SPA can be implemented in $\mathcal{O}(r\operatorname{nnz}(X))$ operations (essentially, $r$ matrix-vector products). In particular, $X-ve^{\top}$ should not be computed explicitely as $v$ could be dense. See https://gitlab.com/ngillis/robustSPA for a MATLAB implementation.
@@ 8 Numerical experiments
We do not include MVE-SPA as it performs similarly as SPA2, as reported, e.g., in [15, Chapter 7]. The code used to run these experiments is available from https://gitlab.com/ngillis/robustSPA. We do not report computational times as all algorithms are extremely fast, although FAW is about $r$ times slower than SPA, T-SPA and TL-SPA, while SPA2 and TL-SPA2 are about two times slower. (It takes about 10 seconds on a standard laptop to run the experiments described below, running the 6 algorithms on 6120 matrices.)
@@
We use the same synthetic data sets as in  [15, Chapter 7.4.6.1]: $X=WH+N$ where
  • The entries of $W\in\mathbb{R}^{m\times r}$ are generated uniformly at random in the interval $[0,1]$, W = rand(m,r). In this case, $W$ is well-conditioned with high probability. To make $W$ ill-conditioned, we compute the compact SVD of $W=U\Sigma V^{\top}$, then update $W\leftarrow U\Sigma^{\prime}V^{\top}$ where $\Sigma^{\prime}$ is a diagonal matrix whose entries are log-spaced between $10^{-6}$ and 1, hence $\frac{\sigma_{1}(W)}{\sigma_{r}(W)}=10^{6}$ and $\mathcal{K}(W)$ will be smaller, of the order of $5\cdot 10^{5}$; see Table 3.
@@
For each experimental settings, we test 51 different noise levels $\delta$ and, for each noise level, we report the average accuracy (which is the proportion of the indices of the columns of $W$ correctly recovered by the corresponding algorithm) over 30 randomly generated matrices.
Exp. 1: Well-condition Dirichlet experiment
Exp. 2: Ill-condition Dirichlet experiment
Figure 7: Accuracy vs. noise level for Dirichlet experiments.
Exp. 3: Well-condition middle point experiment
Exp. 4: Rank-deficient middle point experiment
Figure 8: Accuracy vs. noise level for the middle point experiments.
Table 2 reports the robustness of each algorithm in each experimental setting. The robustness is defined as the largest noise level for which the corresponding algorithm recovered exactly all indices for all matrices.
@@
  • For the well-conditioned Dirichlet experiment (Exp. 1), all SPA variants perform similarly and have the same robustness. The main reason is that the conditioning of $W$ and its translated variant, $W_{t}$, are close to each other. This is a similar observation as in [15, Fig 7.6].
@@
  • For the well-conditioned middle point experiment (Exp. 3), there is a clear hierarchy: SPA and T-SPA perform similarly (since the matrix is well-conditioned, translation does not bring much improvements), FAW is more robust than TL-SPA but its accuracy decays more rapidly, SPA2 outperforms FAW (as already noted in [15, Fig 7.7]), while TL-SPA2 and SPA2 perform similarly although TL-SPA2 is more robust (see Table 2).
@@
To further validate this observation, Table 3 reports the conditioning of the matrix $W$ after the preprocessing performed by the different algorithms, namely:
  • For SPA: $\mathcal{K}(W)$, since SPA applies directly to the data.
  • For T-SPA and FAW: $\frac{K(W_{t})}{\sigma_{r-1}(W_{t})}$, where $W_{t}$ is the matrix $W$ translated by the column of $X$ with the largest $\ell_{2}$ norm; see Section 2.1.
  • For TL-SPA: $\mathcal{K}(W_{\ell})$, where $W_{\ell}=\begin{pmatrix}W-ve^{\top}\\ c\;e^{\top}\end{pmatrix}$ with $v=Xe/n$ and $c$ given by (7).
  • For SPA2: $\mathcal{K}\left((X(:,\mathcal{J}))^{\dagger}W\right)$, where $\mathcal{J}$ is obtained by SPA.
  • For TL-SPA2: $\mathcal{K}\left((\tilde{X}(:,\mathcal{J}))^{\dagger}W_{\ell}\right)$, where $\mathcal{J}$ is obtained by TL-SPA, and $\tilde{X}=\begin{pmatrix}X-ve^{\top}\\ c\;e^{\top}\end{pmatrix}$ is the translated and lifted version of $X$.
We observe the following:
  • Except for Exp. 2 and $\delta=0.7$, it is always TL-SPA or TL-SPA2 that lead to the lowest condition numbers.
  • For Exp. 1, where $W$ is well conditioned, the variants using the left-inverse preconditioning (that is, TL-SPA2, and to a lesser extent SPA2) lead to high condition numbers for high noise levels (for $\delta\geq 0.46$). However, this is not really worrisome since at these noise levels, the performance of all algorithms degrade significantly (robustness is lost for $\delta<0.136$, see Table 2).
  • Exp. 2, where $W$ is ill conditioned, TL-SPA2 is very effective at reducing the conditioning, especially for low noise levels, which explains its robustness results from Table 2.
  • Exp. 4, where $W$ is rank deficient, TL-SPA2 is very effective in reducing the conditioning, when compared to T-SPA and TL-SPA, again confirming the results from Table 2.
Table 3: Average condition number, and standard deviation, of the matrix $W$ after preprocessing using the different algorithms, for the 30 generated matrices (same settings as in Table 2). The first column reports the noise level $\delta$. The best result is highlighted in bold.
Exp. 1 SPA T-SPA & FAW TL-SPA SPA2 TL-SPA2
0.10 $3.7\pm 0.4$ $2.7\pm 0.3$ $4.5\pm 1.4$ $\mathbf{1.2}\pm 0.0$ $\mathbf{1.2}\pm 0.0$
0.28 $3.9\pm 0.4$ $2.8\pm 0.4$ $1.9\pm 0.2$ $1.8\pm 0.2$ $\mathbf{1.7}\pm 0.2$
0.46 $3.7\pm 0.4$ $2.9\pm 0.4$ $\mathbf{2.0}\pm 0.2$ $23.6\pm 25.8$ $22.6\pm 52.8$
0.64 $3.7\pm 0.5$ $3.1\pm 0.4$ $\mathbf{2.1}\pm 0.2$ $75.4\pm 241.1$ $212.4\pm 775.5$
0.82 $3.8\pm 0.3$ $3.6\pm 0.4$ $\mathbf{2.5}\pm 0.2$ $58.1\pm 78.8$ $135.5\pm 501.8$
1 $3.8\pm 0.3$ $4.0\pm 0.5$ $\mathbf{2.6}\pm 0.2$ $49.1\pm 60.1$ $283.6\pm 621.0$
Exp. 2 SPA T-SPA & FAW TL-SPA SPA2 TL-SPA2
$10^{-6}$ $54.7\pm 5.55\cdot 10^{4}$ $20.4\pm 1.93\cdot 10^{4}$ $11.9\pm 1.20\cdot 10^{4}$ ${1.1}\pm 0.0$ $\mathbf{1.0}\pm 0.0$
[…]
0.003 $58.5\pm 7.81\cdot 10^{4}$ $48.4\pm 4.63\cdot 10^{3}$ $12.8\pm 1.74\cdot 10^{4}$ ${1.13}\pm 1.51\cdot 10^{4}$ $\mathbf{4.57}\pm 8.37\cdot 10^{3}$
0.047 $58.5\pm 7.39\cdot 10^{4}$ ${49.6}\pm 4.59\cdot 10^{3}$ $12.7\pm 1.56\cdot 10^{4}$ $8.01\pm 30.9\cdot 10^{5}$ $\mathbf{2.19}\pm 3.13\cdot 10^{4}$
0.7 $57.4\pm 6.06\cdot 10^{4}$ $\mathbf{49.1}\pm 4.68\cdot 10^{3}$ $12.7\pm 1.26\cdot 10^{4}$ $2.21\pm 5.67\cdot 10^{6}$ $4.31\pm 5.94\cdot 10^{5}$
Exp. 3 SPA T-SPA & FAW TL-SPA SPA2 TL-SPA2
0.10 $3.9\pm 0.4$ $2.7\pm 0.3$ $2.2\pm 0.2$ $\mathbf{1.0}\pm 0.0$ $\mathbf{1.0}\pm 0.0$
0.16 $3.7\pm 0.4$ $2.6\pm 0.3$ $2.2\pm 0.2$ $\mathbf{1.0}\pm 0.0$ $\mathbf{1.0}\pm 0.0$
0.25 $3.7\pm 0.4$ $2.6\pm 0.3$ $2.2\pm 0.2$ $1.6\pm 0.7$ $\mathbf{1.2}\pm 0.5$
0.40 $3.7\pm 0.4$ $2.6\pm 0.3$ $\mathbf{2.2}\pm 0.2$ $2.7\pm 0.2$ $2.7\pm 0.2$
0.63 $3.8\pm 0.6$ $3.0\pm 0.4$ $\mathbf{2.3}\pm 0.3$ $3.2\pm 0.4$ $3.1\pm 0.3$
1 $3.9\pm 0.4$ $3.3\pm 0.4$ $\mathbf{2.3}\pm 0.2$ $3.5\pm 0.4$ $3.6\pm 0.4$
Exp. 4 SPA T-SPA & FAW TL-SPA SPA2 TL-SPA2
0.01 $\infty$ $72.0\pm 95.5$ $57.7\pm 73.8$infty $\mathbf{1.0}\pm 0.0$
0.025 $\infty$ $53.1\pm 72.0$ $42.6\pm 57.6$infty $\mathbf{1.0}\pm 0.0$
0.063 $\infty$ $7.65\pm 38.3\cdot 10^{3}$ $6.51\pm 32.8\cdot 10^{3}$infty $\mathbf{1.2}\pm 0.5$
0.16 $\infty$ $407\pm 1610$ $324\pm 1252$infty $\mathbf{1.8}\pm 0.7$
0.40 $\infty$ $56.5\pm 94.6$ $41.4\pm 72.1$infty $\mathbf{2.9}\pm 0.4$
1 $\infty$ $195\pm 600$ $129\pm 416$infty $\mathbf{3.5}\pm 0.4$
Remark 5 (Sensitivity of TL-SPA to the lift parameter).
In Section 7, we proposed a lift parameter for TL-SPA based on Lemma 5; see (7). This is a heuristic choice with no theoretical guarantees since $W$ is unknown in practice. However, we have observed that TL-SPA is not too sensitive to the choice of this parameter. We have rerun TL-SPA with the lift parameter $c$ defined in (7) multiplied by a factor $\alpha\in\{0.1,0.5,2,10\}$. Table 4 reports the average condition number, $\mathcal{K}(W_{\ell})$, for all experiments. Except when $\alpha=10$ that leads to significantly worse results, the other condition numbers are very close to each other.
Table 4: Average conditioning $\mathcal{K}(W_{\ell})$ using different shifts $c$ defined in (7) multiplied by a factor $\alpha\in\{0.1,0.5,2,10\}$. Same settings as in Table 3 but the results are averaged over all noise levels. The best result is highlighted in bold.
$\alpha=1/10$ $\alpha=1/2$ $\alpha=1$ $\alpha=2$ $\alpha=10$
Exp. 1 $6.3\pm 6.4$ $3.0\pm 2.6$ $\mathbf{2.6}\pm 1.1$ $3.0\pm 0.9$ $10.6\pm 6.2$
Exp. 2 $\mathbf{12.4}\pm 1.60\cdot 10^{4}$ $\mathbf{12.4}\pm 1.58\cdot 10^{4}$ $12.5\pm 1.57\cdot 10^{4}$ $13.0\pm 1.54\cdot 10^{4}$ $22.9\pm 2.44\cdot 10^{4}$
Exp. 3 $4.2\pm 0.3$ $\mathbf{1.9}\pm 0.2$ $2.2\pm 0.2$ $3.2\pm 0.4$ $13.6\pm 1.6$
Exp. 4 $\mathbf{1.03}\pm 11.8\cdot 10^{3}$ $1.07\pm 12.2\cdot 10^{3}$ $1.18\pm 13.4\cdot 10^{3}$ $1.56\pm 17.6\cdot 10^{3}$ $5.92\pm 66.4\cdot 10^{3}$
@@ 9 Conclusion and further work
In this paper, we revisited the robustness to noise of SPA and variants. We improved previous error bounds for the first step of SPA and when $r=2$ (Theorems 3 and 4), and for the first two steps and when $r=3$ for the translated variant of SPA, T-SPA (Theorems 5 and 6). We then proved that, for $r\geq 3$, the original robustness result of SPA [19, Theorem 3] is tight (Theorem 7), an important open question in the literature [15, p. 231], using a well-constructed family of matrices that achieve the worst-case error bound. We also proved the tightness of the error bounds for two preconditioned variants of SPA, namely SPA preconditioned with itself (SPA2, Theorem 9) and SPA preconditioned with minimum-volume ellipsoid (MVE-SPA, Theorem 11). Finally, we proposed a new SPA variant that uses a translation and lifting of the data points, namely TL-SPA, and that allows us to improve the robustness of SPA by reducing the condition number of $W$. We illustrated this on numerical experiments.
@@
Acknowledgments
We are grateful to the anonymous reviewers who carefully read the manuscript, their feedback helped us improve our paper.